Score: 0 of 1 pt For the provided sample mean, sample size, and population standard deviation, complete parts (a) through (c) below. x#21,n-100, σ 2 Find a 95% confidence interval for the population mean. The 95% confidence interval is from | | to (Round to two decimal places as needed.) Enter your answer in the edit fields and then click Check Answer. parts remaining
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 21

Population standard deviation = \sigma = 2

Sample size = n = 100

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * ( 2 / \sqrt 100)

= 0.39

At 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

21 - 0.39 < \mu < 21 + 0.39

20.61 < \mu < 21.39

The 95% confidence interval is from (20.61 to 21.39)